# The Kinematic Properties of TŻO Candidate HV 11417 with Gaia DR3

### v1

@@ 1 Introduction
Thorne-Żytkow Objects (TŻOs), first theorized by Thorne & Żytkow (1975, 1977), are a theoretical and intriguing avenue of binary evolution, wherein a neutron star (NS) is engulfed by the envelope of a companion red supergiant (RSG) and merges with its core, forming a single object. Significant work has gone into formalizing the physics of their formation (Taam et al., 1978; Ray et al., 1987; Podsiadlowski et al., 1995; Fryer et al., 1996; Everson et al., 2024; Hutchinson-Smith et al., 2024; Nathaniel et al., 2025; Williams et al., 2025), evolution (Cannon et al., 1992; Cannon, 1993; Biehle, 1994; Podsiadlowski et al., 1995; Farmer et al., 2023), final fates (Podsiadlowski et al., 1995; Liu et al., 2015; Moriya, 2018; Moriya & Blinnikov, 2021), and observational properties – see O’Grady et al. (2026) for a review.
To date, the best studied TŻO candidate is HV 2112 (Levesque et al., 2014), a cool, luminous star in the Small Magellanic Cloud (SMC), though its true nature has been debated (Tout et al., 2014; Maccarone & de Mink, 2016; Beasor et al., 2018; McMillan & Church, 2018; O’Grady et al., 2020, 2023; Farmer et al., 2023). A second TŻO candidate, HV 11417, also located in the SMC, was identified as a candidate by Beasor et al. (2018). The candidacy of this star has also been debated, both from the perspective of new TŻO models (Farmer et al., 2023), but also because its membership within the SMC was brought into question by O’Grady et al. (2020). Using Gaia DR2 (Gaia Collaboration et al., 2018a) kinematics, O’Grady et al. (2020) found that the proper motion of HV 11417 was marginally inconsistent with the proper motions of SMC stars, suggesting it was instead a foreground star and thus far too dim to be a TŻO. However, measurement uncertainties within Gaia DR2, especially for the of HV 11417 (), were very large.
@@ 4.2 Implications for classification
If HV 11417 is a TŻO, its kinematic properties may reflect its formation history, as described in §2.1 of O’Grady et al. (2023). Leonard et al. (1994) predicts the median velocity for TŻOs formed by the natal kick launching the NS into the secondary to be km/s. Assuming generally isotropic velocities, this corresponds to a local transverse velocity of v km/s. Alternatively, for TŻOs formed from the common envelope spiral-in after a HMXB phase (Taam et al. 1978, but see also Papish et al. 2015; Hutchinson-Smith et al. 2024; Everson et al. 2024) could have velocities in the range of 10-80 km/s (van den Heuvel et al., 2000), corresponding to v 8-65 km/s.
The local transverse velocity of HV 11417 of km/s is thus consistent with either of these formation scenarios. Combined with the luminosity (log(L/L 4.92) and elemental abundance anomalies uncovered by Beasor et al. (2018) (particularly the enhanced Rb; although Beasor et al. (2018) did not detect enhanced Li, see §5.7.3 of O’Grady et al. (2023) for discussion) the properties of HV 11417 indeed align with predictions from earlier TŻO models (Thorne & Żytkow, 1975, 1977; Cannon et al., 1992; Cannon, 1993).
@@
However, Farmer et al. (2023) recently created new models of TŻOs with MESA, finding several differences compared to the older models of Cannon et al. (1992); Cannon (1993) – though note that the structure of these new models was only computed down to the point (colloquially known as the ‘knee’) where the envelope becomes fully convective. The luminosity, metallicity, and variability period ( 1090 day fundamental mode) of HV 11417 cannot be consistently matched by any of the Farmer et al. (2023) models, thus Farmer et al. (2023) rule HV 11417 out as a TŻO candidate.
@@ Acknowledgments
The author thanks Maria Drout, Katie Breivik, and Adiv Paradise for useful conversations. The author gratefully acknowledges support from the McWilliams Postdoctoral Fellowship in the McWilliams Center for Cosmology and Astrophysics at Carnegie Mellon University.
@@
Software: This work made use of the following software packages: astropy (Astropy Collaboration et al., 2013, 2018, 2022), Jupyter (Perez & Granger, 2007; Kluyver et al., 2016), matplotlib (Hunter, 2007), numpy (Harris et al., 2020), pandas (McKinney, 2010; pandas development team, 2025), pyia (Price-Whelan & Brammer, 2021), python (Van Rossum & Drake, 2009), and scipy (Virtanen et al., 2020; Gommers et al., 2025).
@@
Software citation information aggregated using The Software Citation Station (Wagg & Broekgaarden, 2024; Wagg et al., 2025).

### v2

@@ 1 Introduction
Thorne-Żytkow Objects (TŻOs), first theorized by Thorne and Żytkow (1975, 1977), are a theoretical and intriguing avenue of binary evolution, wherein a neutron star (NS) is engulfed by the envelope of a companion red supergiant (RSG) and merges with its core, forming a single object. Significant work has gone into formalizing the physics of their formation (Taam et al., 1978; Ray et al., 1987; Podsiadlowski et al., 1995; Fryer et al., 1996; Everson et al., 2024; Hutchinson-Smith et al., 2024; Nathaniel et al., 2025; Williams et al., 2025), evolution (Cannon et al., 1992; Cannon, 1993; Biehle, 1994; Podsiadlowski et al., 1995; Farmer et al., 2023), final fates (Podsiadlowski et al., 1995; Liu et al., 2015; Moriya, 2018; Moriya and Blinnikov, 2021), and observational properties – see O’Grady et al. (2026) for a review.
To date, the best studied TŻO candidate is HV 2112 (Levesque et al., 2014), a cool, luminous star in the Small Magellanic Cloud (SMC), though its true nature has been debated (Tout et al., 2014; Maccarone and de Mink, 2016; Beasor et al., 2018; McMillan and Church, 2018; O’Grady et al., 2020, 2023; Farmer et al., 2023). A second TŻO candidate, HV 11417, also located in the SMC, was identified as a candidate by Beasor et al. (2018). The candidacy of this star has also been debated, both from the perspective of new TŻO models (Farmer et al., 2023), but also because its membership within the SMC was brought into question by O’Grady et al. (2020). Using Gaia DR2 (Gaia Collaboration et al., 2018a) kinematics, O’Grady et al. (2020) found that the proper motion of HV 11417 was marginally inconsistent with the proper motions of SMC stars, suggesting it was instead a foreground star and thus far too dim to be a TŻO. However, measurement uncertainties within Gaia DR2, especially for the of HV 11417 (), were very large.
@@ 4.2 Implications for classification
If HV 11417 is a TŻO, its kinematic properties may reflect its formation history, as described in §2.1 of O’Grady et al. (2023). Leonard et al. (1994) predict the median velocity for TŻOs formed by the natal kick launching the NS into the secondary to be km/s. Assuming generally isotropic velocities, this corresponds to a local transverse velocity of v km/s. Alternatively, TŻOs formed from the common envelope spiral-in after a HMXB phase (Taam et al. 1978, but see also Papish et al. 2015; Hutchinson-Smith et al. 2024; Everson et al. 2024) could have velocities in the range of 10-80 km/s (van den Heuvel et al., 2000), corresponding to v 8-65 km/s.
The local transverse velocity of HV 11417 of km/s is thus consistent with either of these formation scenarios, though this alone is not enough to distinguish it from evolved stars unbound through non-TŻO scenarios (see Figure 5 of Nathaniel et al. 2025). Combined with the luminosity (log(L/L 4.92) and elemental abundance anomalies uncovered by Beasor et al. (2018) (particularly the enhanced Rb; although Beasor et al. (2018) did not detect enhanced Li, see §5.7.3 of O’Grady et al. (2023) for discussion) the properties of HV 11417 indeed align with predictions from earlier TŻO models (Thorne and Żytkow, 1975, 1977; Cannon et al., 1992; Cannon, 1993).
@@
However, Farmer et al. (2023) recently created new models of TŻOs with MESA, finding several differences compared to the older models of Cannon et al. (1992); Cannon (1993). The luminosity, metallicity, and variability period ( 1090 day fundamental mode) of HV 11417 cannot be consistently matched by any of the Farmer et al. (2023) models, thus Farmer et al. (2023) rule HV 11417 out as a TŻO candidate. We do note that the structure of the Farmer et al. (2023) models is only computed down to the point where the envelope becomes fully convective (colloquially known as the ‘knee’), and instead inject energy at the base of the convective envelope in lieu of modeling the radiative zone and neutron star atmosphere. This may have consequences for the predicted nucleosynthetic yields of Farmer et al. (2023) if material from below the ‘knee’ can be mixed into the convective envelope. As well, invoking a low NS mass of M⊙ allows the models to reach the luminosity of HV 11417 (see Appendix 3 and Figure A1c of Farmer et al. 2023).
@@ Acknowledgments
The author thanks Maria Drout, Katie Breivik, and Adiv Paradise for useful conversations. The author thanks the anonymous referee for helpful suggestions that have improved this manuscript. The author gratefully acknowledges support from the McWilliams Postdoctoral Fellowship in the McWilliams Center for Cosmology and Astrophysics at Carnegie Mellon University.
@@
Software: This work made use of the following software packages: astropy (Astropy Collaboration et al., 2013, 2018, 2022), Jupyter (Perez and Granger, 2007; Kluyver et al., 2016), matplotlib (Hunter, 2007), numpy (Harris et al., 2020), pandas (McKinney, 2010; pandas development team, 2025), pyia (Price-Whelan and Brammer, 2021), python (Van Rossum and Drake, 2009), and scipy (Virtanen et al., 2020; Gommers et al., 2025).
@@
Software citation information aggregated using The Software Citation Station (Wagg and Broekgaarden, 2024; Wagg et al., 2025).